OpenStudy (xapproachesinfinity):

while I was in the bus today I thought about this probability experiment A bus does 5 stops from the starting point to its destination point. In each stop only one individual rides. We want to know what is the probability that exactly 3 women will ride between the stops. Assume that all events are equally likely. Will assume that each stop there is ½ chance that a ride is either a man or a women.

OpenStudy (xapproachesinfinity):

a crazy problem lol

OpenStudy (anonymous):

You thought about this on a bus? Wow! nice job! :)

OpenStudy (xapproachesinfinity):

yeah sometimes i think about math situations like this while doing something hehee

OpenStudy (anonymous):

haha nice!

OpenStudy (xapproachesinfinity):

^_^ :)

OpenStudy (solomonzelman):

Nice

OpenStudy (solomonzelman):

you are having to assume a bit too much, but that makes it even better.... thinking process in it's peak:D

OpenStudy (xapproachesinfinity):

I needed to do those assumption otherwise the problem would be hard hehe

OpenStudy (jtvatsim):

drawing a "probability tree" might be helpful as a start to visualize the situation

OpenStudy (xapproachesinfinity):

that's a good way to start!

OpenStudy (xapproachesinfinity):

i haven't solve it myself but i know the way the answer i guess, i just posted for fun

OpenStudy (xapproachesinfinity):

to the answer*

OpenStudy (jtvatsim):

and it is very fun! :)

OpenStudy (solomonzelman):

probability tree? I never liked it... not that I like statistics, and not even that I really love math. (Only some designated topics of it)

OpenStudy (xapproachesinfinity):

heheh@solomonZelman the probability in general is fun at least to me... although I'm weak in this topic lol

OpenStudy (jtvatsim):

hmm... you might also be able to use a permutation with repeat argument.

OpenStudy (jtvatsim):

b/c what you are really dealing with is trying to figure out how many ways the situation that exactly 3 women will ride the bus. So, that automatically means you have to have exactly 2 men riding the bus. So you are considering sequences like: WWWMM WWMMW WMWMW etc. How many sequences you can make in this fashion will be directly related to the answer.

OpenStudy (xapproachesinfinity):

hmmm that sounds good! but i guess there is no necessity that the number of that will ride will be 2 only

OpenStudy (xapproachesinfinity):

men*

ganeshie8 (ganeshie8):

looks like a binomial distribution ?\[P(X=3) = \binom{5}{3}(1/2)^3(1/2)^2\]

OpenStudy (xapproachesinfinity):

yes! that's what i was thinking about! actually what made me think about probability today is watching some youtube series lol

OpenStudy (xapproachesinfinity):

https://www.youtube.com/watch?v=8AJPs3gvNlY this the lecturer

ganeshie8 (ganeshie8):

:) jtvatsim's work looks neat, he is actually deriving binomial distribution..

OpenStudy (xapproachesinfinity):

hmm... he/she assumed that only 2 men will ride for the rest of remaining rides

ganeshie8 (ganeshie8):

``` WWWMM WWMMW WMWMW etc. ``` we are picking 3 women from 5 ppl so there will be exactly 5C3 sequences possible

ganeshie8 (ganeshie8):

and thats a correct assumption right ?

OpenStudy (xapproachesinfinity):

hmm i wasn't thinking about that. he/she still need to multiply by the probability of a women riding and the remaining probability

ganeshie8 (ganeshie8):

Ofcourse yes, consider one sequence : `WWWMM` P(W) = 1/2 P(M) = 1/2 so probability for this sequence = 1/2*1/2*1/2*1/2*1/2 = (1/2)^3 * (1/2)^2

ganeshie8 (ganeshie8):

Notice we are accounting for remaining two stops also ^

OpenStudy (misty1212):

why isn't it \[\frac{\binom{5}{2}}{2^5}\]

OpenStudy (xapproachesinfinity):

oh yes, you are right!

ganeshie8 (ganeshie8):

isnt it same as \(\binom{5}{3} \frac{1}{2^5}\) @misty1212

OpenStudy (xapproachesinfinity):

would that be the same, even i take out exactly 3 women.

OpenStudy (xapproachesinfinity):

5 choose 3 and 5 choose 2 are the same @misty1212

OpenStudy (misty1212):

yes of course

OpenStudy (xapproachesinfinity):

then it is the same answer

OpenStudy (misty1212):

you can choose 3 women or choose 2 men

OpenStudy (xapproachesinfinity):

i might have set the problem the wrong way! my concern is choosing 3 women without the assumption that we choose exactly 3 does that make sense lol

ganeshie8 (ganeshie8):

you want "atleast" 3 women ?

OpenStudy (xapproachesinfinity):

yes at least!

ganeshie8 (ganeshie8):

Easy, simply add up remaining two probabilities also : \[ P(X\ge 3) ~~ =~~ P(X=3) + P(X=4) + P(X=5)\]

OpenStudy (xapproachesinfinity):

hmm i see... what the reasoning behind it

OpenStudy (xapproachesinfinity):

is*

OpenStudy (xapproachesinfinity):

this is the same as P(AUBUC) where A,B,C are disjoint

ganeshie8 (ganeshie8):

I think so.. when the events are independent, the probability of success in each event would be same and you can use binomial distribution : \[\large P(X=k) = \binom{n}{k}p^k(1-p)^{n-k}\] where \(p\) = probability of success \(k\) = number of successes

OpenStudy (xapproachesinfinity):

but we don't have independent. surely one event affect the the other from occurring?

ganeshie8 (ganeshie8):

how does next person know what happened previously ?

ganeshie8 (ganeshie8):

he has no clue about what happened in the last event its not like he will have some mysterious knowledge to think : "Oh last event was a woman, so i better not get into the bus this time. let me try the next bus.." right ?

ganeshie8 (ganeshie8):

probability of woman in each event is same : 1/2 it wont change

OpenStudy (xapproachesinfinity):

hmmm makes sense! i agree

ganeshie8 (ganeshie8):

this distribution calc is pretty awesome http://spark.rstudio.com/minebocek/dist_calc/

ganeshie8 (ganeshie8):

you can choose the type of distribution from drop down menu and play with the sliders..

OpenStudy (xapproachesinfinity):

yeah, I'm trying to restudy this stuff i'm watching NJ wildberger as well, but he doesn't explain everything in his statistic series lol

ganeshie8 (ganeshie8):

Ahh that video looks a lot advanced to me.. i only knw the school level stats. idk much about markov chains and other stuff but i remember working on them in linear algebra..

OpenStudy (xapproachesinfinity):

hhh i didn't reach that part yet lol @ganeshie8 thanks, you helped clear a lot^_^

ganeshie8 (ganeshie8):

np :) the distribtion chart makes it easy to see the pverall required probaility for P(X>=3) is 0.5 : http://i.gyazo.com/e5d41ab87baa7d172d5ac75a3c58b477.png

ganeshie8 (ganeshie8):

as usual the area of all bins adds up to 1 and the area of shaded region represents the required probability

OpenStudy (xapproachesinfinity):

i tried it http://prntscr.com/5onc6s

OpenStudy (xapproachesinfinity):

it gives the probability below the chart

OpenStudy (xapproachesinfinity):

very good tool :)

ganeshie8 (ganeshie8):

we can choose the parameters using sliders on left hand side fix \(n =5\), \(p = 0.5\) and try different values of \(a\) i prefer this tool to working the tedius formulas xD